Freight car loadings over an 18-week period at a busy port are as follows: Week Number Week Number Week Number 1 300 7 365 13 485 2 305 8 395 14 490 3 310 9 425 15 540 4 305 10 420 16 545 5 310 11 450 17 550 6 315 12 480 18 555

a. Determine a linear trend line for expected freight car loadings. (Round your intermediate calculations and final answers to 2 decimal places.) Yˆ = + t

b. Use the above trend equation to predict expected loadings for Weeks 20 & 21. (Round your final answers to 2 decimal places.) The forecasted demand for Week 20 is and for Week 21 is .

c. The manager intends to install new equipment when the volume exceeds 880 loadings per week. Assuming the current trend continues, in which week (at the earliest) should the loading volume reach that level? (Use the rounded answers, as required, from any previous part of this problem. Do not round any other intermediate calculations. Round your final answer to 2 decimal places.) It should reach 880 loadings in Week .
Answer #1

a.

Using excel for performing regression with freight car loadings as the dependent variable and week as independent variable.

In excel ,go to data tab-->data analysis and select regression

In the regression dialog select the freight car loadings values as Y range and week values as x range and click OK.

We get the regression output with coefficients of week and intercept .

The trend equation is Expected freight car loadings =Intercept+coefficient of week * week

Expected freight car loadings=250.59+17.75 t

f $F$7+B21*$F$8 C1 A C E F G H K M N Freight Car loadings Trend forecast 300 Week 1 Regression 2 268.33 305 286.08 Input OK 3

b. Forecasted demand for week 20=250.59+17.75*20=605.49

  Forecasted demand for week 20=250.59+17.75*21=623.24

c. In excel column B is Week and D is trend forecast

Expected freight car loading exceeds 880 in week 36

f$F$7+B37*$F$8 D37 A C D E E 20 21 605,49 22 21 623,24 23 22 640.98 24 23 658.73 676.47 25 24 694.22 26 25 27 26 711.96 28 27